# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v3

@@ IV Discretely coupled FitzHugh-Nagumo equations
F Goetze and Lai (2022); Vaidyanathan (2015); Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Stefanescu and Jirsa (2008); Li et al. (2021); Binczak et al. (2006); Yanagita et al. (2005) F T Wang et al. (2009); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Toral et al. (2003b); Yu et al. (2021); Brandstetter et al. (2010); Panchuk et al. (2013); Ciszak et al. (2003); Schöll et al. (2009); Saha and Feudel (2017) T F Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Keener (1988); Lindner (2004); Brown et al. (2004); Neiman et al. (1999); Heinrich et al. (2010); Perc (2005a); Takembo et al. (2019) F Lei and Yang (2021); Schöll (2016a); Omelchenko et al. (2013); Nikitin et al. (2019); Semenova et al. (2016); Schülen et al. (2022); Schöll (2016b); Omelchenko et al. (2015); Perc (2007); Steur et al. (2009); Zhang and Liao (2017); Chouzouris et al. (2018); Zaks et al. (2005); Vodeneev et al. (2011); Sukhov et al. (2011); Schmidt et al. (2017); García-Ojalvo and Schimansky-Geier (1999); Cartwright (2000); Gerster et al. (2020); Scialla et al. (2021); Ullner et al. (2003b) T Burić and Todorović (2003); Kori et al. (2012); Perlikowski et al. (2010) T Sawicki et al. (2019); Schülen et al. (2019) Discretely coupled ODEs, describing various network configurations, have garnered substantial attention in diverse fields such as neuroscience Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020), electrical systems Heinrich et al. (2010); Zhang and Liao (2017); Li et al. (2021); Binczak et al. (2006); Takembo et al. (2019); Ciszak et al. (2003), and biology Kori et al. (2012); Vodeneev et al. (2011); Sukhov et al. (2011); Cartwright (2000), owing to their relevance and the broad spectrum of topologies and coupling terms they accommodate. This versatility allows for the exploration of various phenomena, enriching the mathematical study of networks Vaidyanathan (2015); Hoff et al. (2014); Lei and Yang (2021); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Lindner (2004); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Semenova et al. (2016); Schülen et al. (2022); Schöll (2016b); Omelchenko et al. (2015); Perc (2007, 2005a); Steur et al. (2009); Zaks et al. (2005); Schmidt et al. (2017); García-Ojalvo and Schimansky-Geier (1999); Gerster et al. (2020); Yanagita et al. (2005).
@@
Network topology, concerning the structure and connectivity of nodes (Fig. 15A), combined with the diversity in coupling terms—from simple functional forms to time-delayed terms—sets the stage for endless modeling possibilities. The classification of networks, as per the literature, often revolves around the network size, the extent of node connections, and the nature of coupling terms, encompassing both functional forms and time delays (Fig. 15B).
@@
The coupled FHN models have primarily been employed to investigate synchronization phenomena Goetze and Lai (2022); Vaidyanathan (2015); Toral et al. (2003a); Kori et al. (2012); Ambrosio and Aziz-Alaoui (2012); Lindner (2004); Brown et al. (2004); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Perc (2007, 2005a); Stefanescu and Jirsa (2008); Steur et al. (2009); Zhang and Liao (2017); Chouzouris et al. (2018); Zaks et al. (2005); Li et al. (2021); Schmidt et al. (2017); Binczak et al. (2006); Cartwright (2000); Gerster et al. (2020); Yanagita et al. (2005); Wang et al. (2009); Toral et al. (2003b); Yu et al. (2021); Brandstetter et al. (2010); Nikitin et al. (2019); Sawicki et al. (2019); Panchuk et al. (2013); Schülen et al. (2019); Ciszak et al. (2003); Schöll et al. (2009); Perlikowski et al. (2010); Saha and Feudel (2017), including the study of chimera states Schöll (2016a); Omelchenko et al. (2013); Nikitin et al. (2019); Sawicki et al. (2019); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Chouzouris et al. (2018); Schmidt et al. (2017). Additionally, they’ve been used to explore stability properties Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Heinrich et al. (2010); Schülen et al. (2022); Zhang and Liao (2017); Vodeneev et al. (2011); Sukhov et al. (2011); Li et al. (2021); Yanagita et al. (2005); Wang et al. (2009); Burić and Todorović (2003); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Perlikowski et al. (2010); Saha and Feudel (2017), traveling waves Keener (1988); Vodeneev et al. (2011); Sukhov et al. (2011); McKean Jr (1970); Sawicki et al. (2019), pattern formation Ambrosio and Aziz-Alaoui (2012); Lindner (2004); García-Ojalvo and Schimansky-Geier (1999); Lei and Yang (2021); Takembo et al. (2019); Perlikowski et al. (2010); Neiman et al. (1999); Schülen et al. (2022); Perc (2007); Schmidt et al. (2017); Sawicki et al. (2019), and the influence of noise on dynamics Toral et al. (2003a); Yu et al. (2021); Lindner (2004); Neiman et al. (1999); Brandstetter et al. (2010); Semenova et al. (2016); Perc (2007, 2005a); Zaks et al. (2005); García-Ojalvo and Schimansky-Geier (1999).
@@
Given the vast array of possibilities with coupled ODEs, this review will focus on selected popular topics within the FHN model context. In Subsection IV.1, we discuss the transition from continuous systems to discrete coupled ODEs, highlighting the evolution towards discretely networked configurations. Subsection IV.2 looks into one of the most basic forms of coupling—the functional difference—and focuses on a two-node system Hoff et al. (2014); Campbell and Waite (2001). This part provides insights into stability analysis and synchronization dynamics, employing the simplest coupling scenarios frequently found in literature Vaidyanathan (2015); Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Keener (1988); Lindner (2004); Neiman et al. (1999); Heinrich et al. (2010); Perc (2005a); Stefanescu and Jirsa (2008); Yanagita et al. (2005). Finally, subsection IV.3 aims to elucidate chimera states, a key area of interest within the FHN model studies, by replicating and discussing the findings of Omelchenko et al. Omelchenko et al. (2013, 2015). This section underscores the significance of chimeras in the study of complex network dynamics.
Figure 16: Transition from continuous to discrete systems in wave propagation A. A continuous system with periodic boundary conditions (left) is contrasted against its discrete counterpart (right). The presence or absence of traveling waves depends on whether the excited region propagates. B. Quantifying wave propagation when transitioning from a continuous system (left) to discrete systems (right), under parameters set to favor the active state ($a=-0.1$, $b=2$, $\varepsilon=0.01$). As the number of coupled nodes increases (32, 128, and 256 nodes examined), the region supporting wave propagation expands, progressively mirroring the continuous system’s behavior. The activation percentages are delineated by the proportion of initially active nodes.
@@ IV.1 From continuously to discretely coupled FHN systems
Alternatively, networks might adopt various topologies based on functional differences, influencing the coupling dynamics Vaidyanathan (2015); Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Keener (1988); Lindner (2004); Neiman et al. (1999); Heinrich et al. (2010); Perc (2005a); Stefanescu and Jirsa (2008); Yanagita et al. (2005). A common form of expressing this coupling is:
$\begin{array}[]{ @{} l @{} }{u_{i}}_{t}=f(u_{i},v_{i})+\eta(-2u_{i}+u_{i+1}+u_{i-1})\\[3.0pt] {v_{i}}_{t}=\varepsilon g(u_{i},v_{i})+\eta(-2v_{i}+v_{i+1}+v_{i-1})\end{array},$ (31)
@@
An intriguing aspect is how wave propagation changes with the number of FHN units or discretization step. Keener’s work on simulating traveling waves in a cardiac model using discrete FHN nodes serves as a notable example Keener (1988), alongside other diverse applications Keener (1988); Vodeneev et al. (2011); Sukhov et al. (2011); McKean Jr (1970); Sawicki et al. (2019). Figure 16 shows the division in a continuous system (with a high node count) between regions exhibiting traveling waves and those that do not, based on the diffusion coefficient $D$ and the size of the initially excited region (high). This numerical experiment is repeated for discrete node counts of $N=(32,128,256)$. The results clarify that as the number of FHN nodes increases, the threshold for wave propagation converges to that of the continuous system, illustrating the transition between discrete and continuous regimes.
@@ IV.2 A unidirectional functional difference between two coupled FHN models
A significant portion of the literature employing functional coupling consists of functional differences, i.e., coupled ODEs where the differential equation for the $i$-th node is proportional to $\sum_{j}(u_{i}-u_{j})$. Here, we explore the dynamics of a simple two-node system through general functional forms and examine the specific scenario of unidirectional coupling, similar to the studies by Hoff et al. Hoff et al. (2014) and Campbell et al. Campbell and Waite (2001). Campbell et al. assumed no coupling from node 2 to node 1.
@@ IV.2.1 Stationary solutions of two coupled nodes
For a two-node system (Fig. 17A), the stability equations are derived considering general functions $p$, $q$, $r$, and $w$. Focusing on functional differences, we model the scenario where node 2 influences node 1 without reciprocal interaction, embodying the ’slave and master’ dynamic prevalent in neuroscience Toral et al. (2003a). The system equations are given as:
$\begin{array}[]{ @{} l @{} }{u_{1}}_{t}=-u_{1}^{3}+u_{1}-v_{1}+p(u_{1},v_{1},u_{2},v_{2})\\[3.0pt] {v_{1}}_{t}=\varepsilon(u_{1}-bv_{1}+a)+q(u_{1},v_{1},u_{2},v_{2})\\[3.0pt] {u_{2}}_{t}=-u_{2}^{3}+u_{2}-v_{2}+r(u_{1},v_{1},u_{2},v_{2})\\[3.0pt] {v_{2}}_{t}=\varepsilon(u_{2}-bv_{2}+a)+w(u_{1},v_{1},u_{2},v_{2})\end{array},$ (32)
@@
Analyzing the nullclines and fixed points reveals the potential for complex dynamics, extending to multiple stable states under simple unidirectional coupling. Such analysis hinges on the specific functional forms involved, often leading to equations that defy analytical solutions. Thus, we propose general solutions of the form $(u_{1}^{*},v_{1}^{*},u_{2}^{*},v_{2}^{*})$. In our case, although the equations are not analytically solvable, we derive certain simplified relational expressions:
$v_{1}^{*}=\frac{u_{1}^{*}+a}{b},$ $b{u_{1}^{*}}^{3}+(1-b-b\gamma)u_{1}^{*}+b\gamma u_{2}^{*},$ (33)
$v_{2}^{*}=\frac{u_{2}^{*}+a}{b},$ $b{u_{2}^{*}}^{3}+(1-b)u_{2}^{*}+a=0,$ (34)
where the second equation portrays node 2 as operating independently, similar to the scenario without coupling. Conversely, the dynamics of node 1 now exhibit a dependency on $u_{2}$, which could result in mono- or tri-stable behaviors, potentially leading to as many as nine stable fixed points. This elementary derivation highlights how the model’s complexity increases as the power of $3^{N}$, where $N$ is the number of nodes, assuming a simplistic unidirectional coupling scheme and gauging complexity by the number of fixed points.
@@ IV.2.2 Linear stability analysis
Continuing with our analytical approach, we now explore the stability of fixed points by considering small perturbations in the form of $(u_{1},v_{1},u_{2},v_{2})=(u_{1}^{*},v_{1}^{*},u_{2}^{*},v_{2}^{*})+\epsilon(\xi_{u_{1}},\xi_{v_{1}},\xi_{u_{2}},\xi_{v_{2}})e^{\sigma t}+c.c.$, where $\epsilon$ is significantly small. This leads to an eigenvalue problem described by:
$(J-\sigma I_{4\times 4})\begin{pmatrix}\xi_{u_{1}}\\ \xi_{v_{1}}\\ \xi_{u_{2}}\\ \xi_{v_{2}}\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix},$ (35)
@@
$J(u_{1},v_{1},u_{2},v_{2})\equiv\begin{pmatrix}-3u_{1}^{2}+1+\gamma&-1&-\gamma&0\\ \varepsilon&-\varepsilon b&0&0\\ 0&0&-3u_{2}^{2}+1&-1\\ 0&0&\varepsilon&-\varepsilon b\\ \end{pmatrix},$
This setup assumes the functions $p$, $q$, $r$, and $w$ are defined accordingly. The eigenvalues, determined by the Jacobian’s determinant and trace, simplify in this specific context, leading to:
${\rm Det}(J)=-b\varepsilon^{2}(1-3{u_{2}^{*}}^{2})(3b{u_{1}^{*}}^{2}-\gamma b-b+1)-\\ \varepsilon^{2}(-3b{u_{1}^{*}}^{2}+\gamma b+b-1),$ (36)
${\rm Tr}(J)=2-3({u_{1}^{*}}^{2}+{u_{2}^{*}}^{2})-2\varepsilon b+\gamma.$ (37)
These expressions highlight the impact of the chosen coupling functions on system stability.
A Hopf bifurcation arises when ${\rm Tr}(J)=0$, yielding:
${u_{1}^{*}}^{2}+{u_{2}^{*}}^{2}=\frac{2-2\varepsilon b+\gamma}{3}.$ (38)
Notably, setting $\gamma$ to zero retrieves the classic condition for a Hopf bifurcation, $1-\varepsilon b\geq 0$. With $\gamma$, the modified condition becomes $2-2\varepsilon b+\gamma\geq 0$.
This derivation, while generic, applies to a specific example of two coupled FHN equations and is extendable to other coupling types, as illustrated in Fig. 15B. For enthusiasts of numerical bifurcation analysis in ODE systems, we recommend Dhooge et al.’s work Dhooge et al. (2003), which provides a comprehensive bifurcation analysis toolkit Hoff et al. (2014).

### v4

@@ IV Discretely coupled FitzHugh-Nagumo equations
Discretely coupled ODEs, describing various network configurations, have garnered substantial attention in diverse fields such as neuroscience Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020), electrical systems Heinrich et al. (2010); Zhang and Liao (2017); Li et al. (2021); Binczak et al. (2006); Takembo et al. (2019); Ciszak et al. (2003), and biology Kori et al. (2012); Vodeneev et al. (2011); Sukhov et al. (2011); Cartwright (2000), owing to their relevance and the broad spectrum of topologies and coupling terms they accommodate. This versatility allows for the exploration of various phenomena, enriching the mathematical study of networks Vaidyanathan (2015); Hoff et al. (2014); Lei and Yang (2021); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Lindner et al. (2004); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Semenova et al. (2016); Schülen et al. (2022); Schöll (2016b); Omelchenko et al. (2015); Perc (2007, 2005b); Steur et al. (2009); Zaks et al. (2005); Schmidt et al. (2017); García-Ojalvo and Schimansky-Geier (1999); Gerster et al. (2020); Yanagita et al. (2005).
@@
Network topology, concerning the structure and connectivity of nodes (Fig. 15A), combined with the diversity in coupling terms – from simple functional forms to time-delayed terms – sets the stage for endless modeling possibilities. The classification of networks, as per the literature, often revolves around the network size, the extent of node connections, and the nature of coupling terms, encompassing both functional forms and time delays (Fig. 15B).
@@
The coupled FHN models have primarily been employed to investigate synchronization phenomena Goetze and Lai (2022); Vaidyanathan (2015); Toral et al. (2003a); Kori et al. (2012); Ambrosio and Aziz-Alaoui (2012); Lindner et al. (2004); Brown et al. (2004); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Perc (2007, 2005b); Stefanescu and Jirsa (2008); Steur et al. (2009); Zhang and Liao (2017); Chouzouris et al. (2018); Zaks et al. (2005); Li et al. (2021); Schmidt et al. (2017); Binczak et al. (2006); Cartwright (2000); Gerster et al. (2020); Yanagita et al. (2005); Wang et al. (2009); Toral et al. (2003b); Yu et al. (2021); Brandstetter et al. (2010); Nikitin et al. (2019); Sawicki et al. (2019); Panchuk et al. (2013); Schülen et al. (2019); Ciszak et al. (2003); Schöll et al. (2009); Perlikowski et al. (2010); Saha and Feudel (2017), including the study of chimera states Schöll (2016a); Omelchenko et al. (2013); Nikitin et al. (2019); Sawicki et al. (2019); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Chouzouris et al. (2018); Schmidt et al. (2017). Additionally, they have been used to explore stability properties Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Heinrich et al. (2010); Schülen et al. (2022); Zhang and Liao (2017); Vodeneev et al. (2011); Sukhov et al. (2011); Li et al. (2021); Yanagita et al. (2005); Wang et al. (2009); Burić and Todorović (2003); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Perlikowski et al. (2010); Saha and Feudel (2017), traveling waves Keener (1988); Vodeneev et al. (2011); Sukhov et al. (2011); McKean (1970); Sawicki et al. (2019), pattern formation Ambrosio and Aziz-Alaoui (2012); Lindner et al. (2004); García-Ojalvo and Schimansky-Geier (1999); Lei and Yang (2021); Takembo et al. (2019); Perlikowski et al. (2010); Neiman et al. (1999); Schülen et al. (2022); Perc (2007); Schmidt et al. (2017); Sawicki et al. (2019), and the influence of noise on dynamics Toral et al. (2003a); Yu et al. (2021); Lindner et al. (2004); Neiman et al. (1999); Brandstetter et al. (2010); Semenova et al. (2016); Perc (2007, 2005b); Zaks et al. (2005); García-Ojalvo and Schimansky-Geier (1999).
@@
Given the vast array of possibilities with coupled ODEs, this review will focus on selected popular topics within the FHN model context. In IV.1, we discuss the transition from continuous (PDE) systems to discrete coupled ODEs. IV.2 looks into one of the most basic forms of coupling – the functional difference – and focuses on a two-node system Hoff et al. (2014); Campbell and Waite (2001). This part provides insights into stability analysis and synchronization dynamics, employing the simplest coupling scenarios frequently found in literature Vaidyanathan (2015); Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Keener (1988); Lindner et al. (2004); Neiman et al. (1999); Heinrich et al. (2010); Perc (2005b); Stefanescu and Jirsa (2008); Yanagita et al. (2005). Finally, IV.3 discusses so-called chimera states, a key area of interest within the FHN model studies, by replicating and discussing the findings of Omelchenko et al. Omelchenko et al. (2013, 2015). This section underscores the significance of chimeras in the study of complex network dynamics.
@@ IV.1 From continuously to discretely coupled FHN systems
Alternatively, networks might adopt various topologies based on functional differences, influencing the coupling dynamics Vaidyanathan (2015); Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Keener (1988); Lindner et al. (2004); Neiman et al. (1999); Heinrich et al. (2010); Perc (2005b); Stefanescu and Jirsa (2008); Yanagita et al. (2005). A common form of expressing this coupling is:
$\begin{array}[]{ @{} l @{} }{u_{i}}_{t}=f(u_{i},v_{i})+\eta(-2u_{i}+u_{i+1}+u_{i-1}),\vspace{3pt}\\ {v_{i}}_{t}=\varepsilon g(u_{i},v_{i})+\eta(-2v_{i}+v_{i+1}+v_{i-1}),\end{array}$ (30)
@@
An intriguing aspect is how wave propagation changes with the number of FHN units or discretization step. Keener’s work on simulating traveling waves in a cardiac model using discrete FHN nodes serves as a notable example Keener (1988), alongside other diverse applications Keener (1988); Vodeneev et al. (2011); Sukhov et al. (2011); McKean (1970); Sawicki et al. (2019). Fig. 16 shows the division in a continuous system (with a high node count) between regions exhibiting traveling waves and those that do not, based on the diffusion coefficient $D$ and the size of the initially excited region (high). This numerical experiment is repeated for discrete node counts of $N=(32,128,256)$. The results clarify that as the number of FHN nodes increases, the threshold for wave propagation converges to that of the continuous system, illustrating the transition between discrete and continuous regimes.
@@ IV.2 A unidirectional functional difference between two coupled FHN models
A significant portion of the literature employing functional coupling consists of functional differences, i.e., coupled ODEs where the differential equation for the $i$th node is proportional to $\sum_{j}(u_{i}-u_{j})$. Here, we explore the dynamics of a simple two-node system through general functional forms and examine the specific scenario of unidirectional coupling, similar to the studies by Hoff et al. Hoff et al. (2014) and Campbell et al. Campbell and Waite (2001). Campbell et al. assumed no coupling from node 2 to node 1.
Figure 16: Transition from continuous to discrete systems: quantifying wave propagation. A. A continuous system with periodic boundary conditions (left) is contrasted against its discrete counterpart (right). The presence or absence of traveling waves depends on whether the excited region propagates. B. Quantifying wave propagation when transitioning from a continuous system (left) to discrete systems (right), under parameters set to favor the active state ($a=-0.1$, $b=2$, $\varepsilon=0.01$). As the number of coupled nodes increases (32, 128, and 256 nodes examined), the region supporting wave propagation expands, progressively mirroring the continuous system’s behavior. The activation percentages are delineated by the proportion of initially active nodes.
@@ IV.2.1 Stationary solutions of two coupled nodes
For a two-node system (Fig. 17A), the stability equations are derived considering general functions $p$, $q$, $r$, and $w$. Focusing on functional differences, we model the scenario where node 2 influences node 1 without reciprocal interaction, embodying the ‘slave and master’ dynamic prevalent in neuroscience Toral et al. (2003a). The system equations are given as:
$\begin{array}[]{ @{} l @{} }{u_{1}}_{t}=-u_{1}^{3}+u_{1}-v_{1}+p(u_{1},v_{1},u_{2},v_{2}),\vspace{3pt}\\ {v_{1}}_{t}=\varepsilon(u_{1}-bv_{1}+a)+q(u_{1},v_{1},u_{2},v_{2}),\vspace{3pt}\\ {u_{2}}_{t}=-u_{2}^{3}+u_{2}-v_{2}+r(u_{1},v_{1},u_{2},v_{2}),\vspace{3pt}\\ {v_{2}}_{t}=\varepsilon(u_{2}-bv_{2}+a)+w(u_{1},v_{1},u_{2},v_{2}),\end{array}$ (31)
@@
Analyzing the nullclines and fixed points reveals the potential for complex dynamics, extending to multiple stable states under simple unidirectional coupling. Such analysis hinges on the specific functional forms involved, often leading to equations for which there are no analytical solutions. Thus, we propose general solutions of the form $(u_{\text{FP},1},v_{\text{FP},1},u_{\text{FP},2},v_{\text{FP},2})$. In our case, although the equations are not analytically solvable, we derive certain simplified relational expressions:
$\begin{array}[]{ll}v_{\text{FP},1}=\frac{u_{\text{FP},1}+a}{b},&bu_{\text{FP},1}^{3}+(1-b-b\gamma)u_{\text{FP},1}+b\gamma u_{\text{FP},2},\vspace{3pt}\\ v_{\text{FP},2}=\frac{u_{\text{FP},2}+a}{b},&bu_{\text{FP},2}^{3}+(1-b)u_{\text{FP},2}+a=0,\end{array}$ (32)
where the second equation represents node 2 as operating independently, similar to the scenario without coupling. Conversely, the dynamics of node 1 now exhibit a dependency on $u_{2}$, which could result in mono- or tri-stable behaviors, potentially leading to as many as nine stable fixed points. This elementary derivation highlights how the model’s complexity increases as the power of $3^{N}$, where $N$ is the number of nodes, assuming a simplistic unidirectional coupling scheme and gauging complexity by the number of fixed points.
@@ IV.2.2 Linear stability analysis
Continuing with this analytical approach, we now explore the stability of fixed points by considering small perturbations in the form of
$(u_{1},v_{1},u_{2},v_{2})=(u_{\text{FP},1},v_{\text{FP},1},u_{\text{FP},2},v_{\text{FP},2})+\epsilon(\xi_{u_{1}},\xi_{v_{1}},\xi_{u_{2}},\xi_{v_{2}})e^{\sigma t}+c.c.,$
where $\epsilon$ is sufficiently small. This leads to an eigenvalue problem described by:
$(J-\sigma I_{4\times 4})\begin{pmatrix}\xi_{u_{1}}\\ \xi_{v_{1}}\\ \xi_{u_{2}}\\ \xi_{v_{2}}\end{pmatrix}=\begin{pmatrix}0\\ 0\\ 0\\ 0\end{pmatrix},$ (33)
@@
$J\equiv\begin{pmatrix}-3u_{\text{FP},1}^{2}+1+\gamma&-1&-\gamma&0\\ \varepsilon&-\varepsilon b&0&0\\ 0&0&-3u_{\text{FP},2}^{2}+1&-1\\ 0&0&\varepsilon&-\varepsilon b\end{pmatrix}.$
Unlike two-dimensional systems, the eigenvalues cannot be calculated solely using the trace and determinant. Instead, computational or advanced analytical methods are required to solve the characteristic polynomial:
$\text{det}(J-\sigma I_{4\times 4})=\left[\sigma^{2}-(\kappa_{1}-\varepsilon b)\sigma+\varepsilon-\kappa_{1}\varepsilon b\right]\left[\sigma^{2}-(\kappa_{2}-\varepsilon b)\sigma+\varepsilon-\kappa_{2}\varepsilon b\right],$ (34)
where $\kappa_{1}=-3u_{\text{FP},1}^{2}+1+\gamma$ and $\kappa_{2}=-3u_{\text{FP},2}^{2}+1$. Analogous to the zero-trace condition, a Hopf bifurcation occurs when a pair of eigenvalues becomes purely imaginary, while the remaining eigenvalues have negative real parts. Thus, $\kappa_{i}-\varepsilon b=0$ and $01-\kappa_{i}b>$ 0 for $i=1,2$, while $\mathbb{R}[\kappa_{j}-\varepsilon b\pm\sqrt{(\kappa_{j}-\varepsilon b)^{2}-4\varepsilon(1-\kappa_{j}b)}]\leq 0$ for $j\neq i$.
This derivation can be extended to other coupling types, as illustrated in Fig. 15B. In certain scenarios, the symmetric properties of the coupling and the fixed points can simplify the analysis, yielding more straightforward analytical expressions Campbell and Waite (2001). For those interested in numerical bifurcation analysis in ODE systems, we recommend the work of Dhooge et al. Dhooge et al. (2003), which offers a comprehensive bifurcation analysis toolkit Hoff et al. (2014).